OpenStudy (anonymous):

A flagpole broken over by the wind forms a right triangle w/ the ground.if the angle w/c the broken part makes w/ the ground is 50 degree and the distance from the tip of the pole to the foot is 55 feet how tall was the pole?

OpenStudy (anonymous):

 So you need to find p1 and p2 in my drawing. p1 is opposite the 50 degree angle, so it is: tan(50) = p1/55 p2 is the hypotenuse cos(50) = p2/55 height = p1 + p2 = 55(tan(50) + cos(50))